OpenStudy (anonymous):

Evaluate 5 − 4b + c if b = -2 and c = -5.

OpenStudy (anonymous):

Plug in your values: \[5-4b+c=5-(4)(-2) + (-5)= 5--8 -5 =13-5=8\]

OpenStudy (anonymous):

5 − 4b + c 5-4(-2)+(-5) 5+8-5 8